# Driven-dissipative turbulence in exciton-polariton quantum fluids

### v2

@@ I Introduction
Quantum turbulence is a peculiar stochastic phenomenon characterized by the emergence of a spontaneous order from chaotic motion, due to the formation of self-similar spatial structures of vorticity topological defects in a quantum fluid. Contrary to the observations of 3D quantum turbulence in quantum systems, such as superfluid helium [1, 2] and ultra-cold atomic Bose-Einstein condensates [3, 4], the existence of an inverse cascade for 2D quantum turbulence analogous to the one observed in classical fluids is still under strong debate, due to both controversial scaling arguments and very stringent mathematical limits on numerical simulations [5]. Indeed, while several works have numerically confirmed the dynamical clustering of quantized vortices, namely the Onsager vortex clustering, consistently with the inverse energy cascade by means of evaporating heating process [6, 7, 8, 9], other argue against its existence [10]. Today, thanks to recent technological advancements in spectroscopy and semiconductor manufacturing, polariton quantum fluids, with their hybrid light-matter nature, represent an alternative platform for investigating 2D quantum turbulence [11].
@@
The exciton-polariton is a bosonic composite quasiparticle, formed in semiconductor microcavities by the exciton (electron-hole hydrogen-atom-like quasiparticle) strongly coupled to the cavity photon, and undergoes a strong radiative decay because of the finite lifetime of its photonic component. Compared to the long-living atomic systems, this essential feature, in addition to the very small effective mass of the polariton, distinguishes the observed superfluid behavior [12, 13] and Bose-Einstein condensation [14] in polaritonic systems. The latter exhibit a high critical temperature (very small effective mass) and a driven-dissipative nature, which experimentally can be overcome by means of external optical incoherent pumping [15]. The dynamics of topological defects has been extensively investigated. The experimentally estimated healing length value (order of a micrometer) of quantized vortices [16, 17, 18] allows, in principle, the chance to observe an inverse energy cascade over a larger range of magnitude of the wave vector (upper cascade bound ). Moreover, the quantized vortex creation method (in other words, the energy injection method) plays a crucial role in the observation of the energy cascade both theoretically and experimentally, and techniques such as spoon-stirring [19, 20], paddle, or defect grid drag [21, 22] of fluid flowing around stationary defects [23] have been used in the past as a standard vortex injection technique for atomic systems.
@@
The necessity to implement a time-varying potential challenges the implementation of conventional spoon-stirring quantized vortex injection scheme for exciton polariton systems. Currently, among the setups based on the time-varying potential for polariton quantum fluid excitation, one should highlight the single-vortex creation in the rotating-bucket experiments [24, 25, 26]. However, polaritons are suitable for direct wave function imprinting via the quasi-resonant excitation approach with flexible spatial light modulator-based (SLM) phase engineering. This approach was successfully used to excite complex spatial distributions of quantized vortices [27, 28]. It is necessary to find reliable approaches for that type of excitation that are able to reproduce the specificity of conventional spoon stirring in the framework of vortex clustering and spectral energy distribution. In this study, we propose various wave function phase patterns and show that those of tiles of plane waves with randomly directed wave vectors effectively mimic the spoon-stirring method of vortex injection.
@@
The paper is organized as follows. In section II, we describe in details the scheme of the provided numerical simulation of dissipative quantum fluid dynamics within the framework of Gross-Pitaevskii equation. In subsection III.1, the equivalence between imprinting techniques and spoon stirring for quantum turbulence studies is extensively discussed within the framework of spatial distributions, vortex statistics and incompressible kinetic energy (IKE) spectra analysis. Subsection III.2 focuses on the effects of losses on sustainability of quantum turbulence arising from the finite lifetime specificity of polaritons, studying at which extent quantum fluids density depletion effects on IKE spectra and vortex statistics for quantum fluid with losses. Then, in subsection III.3, we investigate the conditions on the existence of reliable quantum turbulence signs in presence of gain/loss fluctuations, modeling the experimental case of a laser operating above and below condensation threshold. The final conclusions and observation are collected in section IV.
@@ II Numerical setup
In the quantum fluids research field, the widely used theoretical model is the mean-field Gross-Pitaevskii equation-based approach, which efficiently describes a Bose-Einstein condensate in equilibrium. Regarding exciton-polaritons, quasiparticles of finite lifetime, such model represents an idealized conservative description and requires modification by adding the specific terms responsible for the gain and loss:
@@
The values for a polariton system are taken as follows: (bare electron mass g), . Eq.(1) is solved by means of third order Adam-Bashforth method for the time integration, with time step ps, total duration ns and a square mesh grid of size m. For the calculation of the Laplace operator, we used the Fourier transform with massive parallelization provided by GPU. The additional -dependent term comes from the hybrid Boltzmann Gross-Pitaevskii theoretical model [30], where the parameter governs the dissipation related to high wave vectors. The value of parameter is predicted to be controlled by the temperature from nearly zero value to that comparable to 0.1. For each simulation, we perform averaging over 16 runs.
@@
In this study, we present and validate two different initial configurations of the condensate wave function and check their validity by comparison with the classical rotating spoon technique. In both cases, we keep the density of the wave function at the level of approximately m-2 which gives healing length m and engineer the phase exclusively. In the first case, we divide the wave function into the square tiles of the size 64 m and fill each of them by randomly oriented plane wave with wave length 8 m (Tile-8) or 4 m (Tile-4). In the second case, we divide wave function into pixels of the size 8 m (Pixel-8) or 4 m (Pixel-4) of the random phase each. Experimentally, this approach can be straightforwardly implemented based on the quasi-resonant pumping and spatial light modulator usage. Total simulation time was ns. During the first (preparatory) part of the simulation (from 0 to ps), polariton lifetime was set sufficiently higher then simulation duration and decay parameter was set to allowing the initial phase pattern evolve to the turbulent state. During the second (analysis) part of the simulation from to , the data for the analysis was collected. For spoon-stirring and tile-imprinting turbulence generation schemes, the density in the beginning of analysis phase was around 120 m-2, while for pixel schemes one obtains around 160 m-2. The confining potential of 480 m in diameter was used. The parameter was switched to zero and lifetime was set depending on the required decay rate. We use the values of polariton lifetime listed in Table 1, which yield the final density 30%, 10%, 3% and 1% of the initial one (for the analysis phase). These values exceed sufficiently the achievable in actual samples, however they allow establishing fundamental limitations for the polariton platform utilization for quantum turbulence studies. The duration of the spoon-stirring numerical experiment was similar, with stirring phase during the first part and analysis during the second part. More simulation details are given in supplemental materials.
0.3 | 0.1 | 0.03 | 0.1 | ||
, ns | 1.89 | 0.99 | 0.65 | 0.5 | 300 |
@@ III.1 Comparison of imprinting techniques and spoon-stirring vortex injection
Fig. 1 illustrates the basic principles of the investigated vortex injection and quantum fluid excitation methods. The panels show various stages of the development of turbulence for the spoon stirring procedure and for the imprinting techniques. The obtained for the spoon-stirring procedure density and phase motifs are typical for that kind of vortex injection method, see panels a) and b). For the imprinting technique Tile-8, we plot (panel c) initial phase pattern, that can be sent to SLM device for quasi-resonant polariton excitation. In panel d), we also plot vorticity map at the very beginning (11 ps) of the free evolution. One sees emergence of vorticity singularities at the tile borders that will further develop into distribution of stable quantized vortices. Finally, in panels e) and f) we plot initial phase pattern for Pixel-8 imprinting technique and the density plot for the well-developed turbulence, respectively. One sees that the phase imprinting techniques are capable of creating the turbulent-like state of quantum fluid. The stochastically created vortex seeds (points with non-zero circulation around) then develop into the well-shaped quantized vortices. The videos of the simulations can be found online, see Supplemental for more details.
@@
Fig. 2 shows the results of vortex detection and the application of the clustering algorithm for the cases of the Tile-8 phase-imprinting technique and for the spoon-stirring procedure at the beginning of the analysis phase (for conservative simulation). Visually, the resulting vortex distributions and their organization into clusters nearly coincide for Tile-8 and spoon-stirring. However, a detailed quantitative analysis of the excitation strategies is necessary from the point of view of cluster statistics and IKE spectra.
@@
To do so, we plot histograms to highlight vortex clusterization for all excitation strategies employed. The left counterparts of the paired columns in panel a) in Fig. 3 show the number of single vortices, vortices in dipoles and vortices in clusters at the beginning of the analysis stage for all excitation techniques employed. One can see that Tile-8 strategy resembles the spoon-stirring one much better than other strategies. Fractions of vortices in single state, dipoles and clusters are sufficiently similar. For Tile-4 and Pixel strategies, one observes substantially lower fraction of vortices in clusters. Panel b) in Fig. 3 gives more details on clusterization for spoon stirring and phase imprinting Tile-8 strategy. Tile-8 gives a slightly higher number of vortices in large clusters whereas the spoon-stirring strategy produces more single vortices, dipoles and small clusters. Typical intervortex distance can be estimated as , which gives m for Tile-8 and spoons-stirring strategies.
@@
IKE spectra calculated via positions of all vortcies and via positions of clustered vortices only (employing Eq. (2)) are depicted in panel a) of Fig. 4 for all excitation strategies. One can see very similar full IKE spectra regardless the strategy employed; the discrepancies are within the values imposed by the difference in the total number of vortices, see Fig. 3. The full IKE spectra are composed of two parts different in characteristic power law. Exponent -3 describes the vortex cores at high wave vectors (greater than ) and exponent -1 dominates at lower wave vectors, being a fingerprint of an uncorrelated gas of vortices with dominating dipoles and single vortices. At the same time, the spectra for clustered vortices demonstrate a visible region of -5/3 power law, as was shown earlier for different excitation strategies [31]. It should be noted that the Tile-8 phase-imprinting excitation strategy demonstrates full and clustered vortex-based IKE spectra very similar to those of the spoon-stirring approach. Among the phase-imprinting techniques, Tile-8 yields the highest clusterization degree comparable to the spoon stirring case. For pixel-type imprinting contribution from clustered vortices is lower, which agrees with the histograms. Panels b) and c) in Fig. 4 show full IKE spectra obtained semi-analytically via Eq. (2) (for full IKE spectra and IKE spectra of clustered vortices) and numerically via Eq. (6) for Tile-8 and spoon stirring cases, respectively. One sees that at low wave vectors (large size scales) full and clustered vortex IKE spectra coincide. This is due to the vanishing global rotational motion from the uncorrelated part of the “vortex gas”. For the wave vectors above , one sees that full IKE spectra have higher magnitude than the clustered vortex IKE spectra proportionally to the fraction of clustered vortices.
@@ III.2 Effects of losses on sustainability of quantum turbulence
In present section, we investigate at which extent quantum fluid density depletion effects on IKE spectra and vortex statistics of quantum fluid with losses (radiative losses in case of polaritons). At this stage of the analysis, it has been clear that Tile-8 imprinting is a valid quantum fluids excitation method to study quantum turbulence, as shown by the analysis of the IKE spectra and vortex statistics. So, we focus on this excitation method. First, we track the time evolution of quantum fluid density and total number of vortices, see Fig. 5 panels a) and b), respectively. One sees that for high lifetime the decay of the total number of vortices is very close to the conservative case. For stronger decay, decrease of total number of vortices is stronger. When the density reaches approximately 20 (or 10% of density in the beginning of analysis phase), an abrupt raise in number of vortices in the automatized detection procedure becomes visible, which corresponds to turbulence vanishment associated with appearance of the noise-like phase pattern. However, for lifetimes higher than 1 ps, turbulence exists for the whole observation period.
@@ III.3 Quantum turbulence in the regimes of fluctuating gain and loss
In the present section, we investigate the possibility of maintaining quantum turbulence in the regime with alternating phases of positive and negative gain. For polaritons, the regime corresponds to the incoherent pump above and below the threshold. We intend to characterize the effects of gain/loss fluctuations, possibly arising either from intentionally modulated lasers or from experimental limitations on the optical laser stability.

### v3

@@ I Introduction
Quantum turbulence is a peculiar stochastic phenomenon characterized by the emergence of a spontaneous order from chaotic motion, due to the formation of self-similar spatial structures of vorticity topological defects in a quantum fluid. Contrary to the observations of 3D quantum turbulence in quantum systems, such as superfluid helium [1, 2] and ultra-cold atomic Bose-Einstein condensates [3, 4], the existence of an inverse cascade for 2D quantum turbulence analogous to the one observed in classical fluids is still under the debate, due to both controversial scaling arguments and very stringent mathematical limits on numerical simulations [5]. Indeed, while several works have numerically confirmed the dynamical clustering of quantized vortices, namely the Onsager vortex clustering, consistently with the inverse energy cascade by means of evaporating heating process [6, 7, 8, 9], other argue against its existence [10]. Today, thanks to recent technological advancements in spectroscopy and semiconductor manufacturing, polariton quantum fluids, with their hybrid light-matter nature, represent an alternative platform for investigating 2D quantum turbulence [11, 12].
@@
The exciton-polariton is a bosonic composite quasiparticle, formed in semiconductor microcavities by the exciton (electron-hole hydrogen-atom-like quasiparticle) strongly coupled to the cavity photon, and undergoes a strong radiative decay because of the finite lifetime of its photonic component. Compared to the long-living atomic systems, this essential feature, in addition to the very small effective mass of the polariton, distinguishes the observed superfluid behavior [13, 14] and Bose-Einstein condensation [15] in polaritonic systems. The latter exhibit a high critical temperature (very small effective mass) and a driven-dissipative nature, which experimentally can be overcome by means of external optical incoherent pumping [16]. The dynamics of topological defects has been extensively investigated. The experimentally estimated healing length value (order of a micrometer) of quantized vortices [17, 18, 19] allows, in principle, the opportunity to observe an inverse energy cascade over a larger range of magnitude of the wave vector (upper cascade bound ). Moreover, the quantized vortex creation method (in other words, the energy injection method) plays a crucial role in the observation of the energy cascade both theoretically and experimentally, and techniques such as spoon-like stirring [20, 21, 22], paddle, or defect grid drag [23, 24] of fluid flowing around stationary defects [25] have been used in the past as a standard vortex injection technique for atomic systems.
@@
The necessity to implement a time-varying potential challenges the implementation of conventional spoon-stirring quantized vortex injection scheme for exciton polariton systems. Currently, among the setups based on the time-varying potential for polariton quantum fluid excitation, one should highlight the creation of one or several vortices using the potentials with rotation-induced chirality [26, 27, 28]. The experimental setups required combination of several lasers with finely controlled frequency difference to make the time-dependent interference-governed gain and potential landscape. The resulting potentials were smooth (and thus less effective in stirring) as a result of the smooth phase variation of the elemental beams. Finally, the long-living (hundreds of picoseconds) exciton reservoir [29, 30] will produce the unnecessary trail. However, polaritons are suitable for direct wave function imprinting via the quasi-resonant excitation approach with flexible spatial light modulator-based (SLM) phase engineering. This approach was successfully used to excite complex spatial distributions of quantized vortices [31, 32]. It is noteworthy that the atomic condensates can also be subjected to the phase imprinting techniques realized by the pulsating potential for e.g. the soliton creation [33]. It is also necessary to find reliable approaches for that type of excitation that are able to reproduce the specificity of conventional spoon stirring in the framework of vortex clustering and spectral energy distribution. In this study, we propose various wave function phase patterns and show that those of tiles of plane waves with randomly directed wave vectors effectively mimic the spoon-stirring method of vortex injection.
@@
The paper is organized as follows. In section II, we describe in detail the scheme of the provided numerical simulation of dissipative quantum fluid dynamics within the framework of Gross-Pitaevskii equation. In subsection III.1, the equivalence between imprinting techniques and spoon stirring for quantum turbulence studies is extensively discussed within the framework of spatial distributions, vortex statistics and incompressible kinetic energy (IKE) spectra analysis. Subsection III.2 focuses on the effects of losses on sustainability of quantum turbulence arising from the finite lifetime specificity of polaritons, studying at which extent quantum fluids density depletion effects on IKE spectra and vortex statistics for quantum fluid with losses. Then, in subsection III.3, we investigate the conditions on the existence of reliable quantum turbulence signs in presence of gain/loss fluctuations, modeling the experimental case of a laser operating above and below condensation threshold. The final conclusions and observation are collected in section IV.
@@ II Numerical setup
In the quantum fluids research field, the widely used theoretical model is the mean-field Gross-Pitaevskii equation-based approach, which efficiently describes a Bose-Einstein condensate in equilibrium. Regarding exciton-polaritons, quasiparticles of finite lifetime, such a model represents an idealized conservative description and requires modification by adding the specific terms responsible for the gain and loss:
@@
The values for a polariton system are taken as follows: (bare electron mass g), . Eq.(7) is solved by means of third order Adam-Bashforth method for the time integration, with time step ps, total duration ns and a square mesh grid of size m. For the calculation of the Laplace operator, we used the Fourier transform with massive parallelization provided by GPU. The additional -dependent term comes from the hybrid Boltzmann Gross-Pitaevskii theoretical model [35], where the parameter governs the dissipation related to high wave vectors. The value of parameter is predicted to be controlled by the temperature from nearly zero value to that comparable to 0.1.
@@
In this study, as one of the principal goals, we present the turbulence generation methods based on direct phase imprinting (governed by the initial wave function ) and compare them with the spoon stirring as a reference. For all methods, the simulation of the total duration ns was divided into the preparatory part and the analysis part. The confining circular hard wall potential of 480 m in diameter was normally used. During the first (preparatory) part of the simulation (from 0 to ps), polariton lifetime was set sufficiently higher than simulation duration and decay parameter was set to to suppress the short wave length density waves that complicate the vortex detection. During the second (analysis) part of the simulation from to , the data for the analysis was collected. The parameter was switched to zero and lifetime was set depending on the required decay rate. We use the values of polariton lifetime listed in Table 1, which yield the final density 30%, 10%, 3% and 1% of the initial one (for the analysis phase). These values exceed sufficiently the achievable in actual samples, however they allow establishing fundamental limitations for the polariton platform utilization for quantum turbulence studies. The duration of the spoon-stirring numerical experiment was similar, with stirring phase during the first part and analysis during the second part. More simulation details, including the videos, are given in the Appendix.
The simulation of spoon stirring was carried out by a 10 m potential propagating with a speed of 1.3 m/ps during the preparatory phase. As for the imprinting methods, the two different families of phase pattern were considered. In the first case, we divide the wave function into the 8x8 grid of the square tiles of size 64 m and fill each of them with a randomly oriented plane wave with wave length 12 m (Tile-12), 8 m (Tile-8) or 4 m (Tile-4). In the second case, we divide wave function into pixels of the size 8 m (Pixel-8) or 4 m (Pixel-4) of the random phase each. In both cases, we keep the density of the wave function at the level of approximately m-2 which gives healing length m. Experimentally, the phase imprinting can be straightforwardly implemented based on the quasi-resonant pumping and spatial light modulator usage as it was done previously for vortex chains [31] and lattices [32]. By the principles, the tile imprinting schemes resemble the orbital angular momentum injection by four arranged in square lasers described in [36] For spoon-stirring and tile-imprinting turbulence generation schemes, the density in the beginning of analysis phase was around 120 m-2, while for pixel schemes one obtains around 160 m-2. Fig. 1 illustrates the basic principles of the investigated vortex injection and turbulent motion excitation methods. Practically, for the imprinting schemes, the preparatory phase (where density waves relaxation and turbulent regime stabilization takes place) could be significantly lower, i.e. around 200 ps. See the simulation videos in the Appendix. This is yet another advantage of the imprinting techniques with respect to the “spoon stirring”.
| 0.3 | 0.1 | 0.03 | ||
| (ns) | 1.89 | 0.99 | 0.65 | 300 |
@@
We averaged the vortex statistics and spectra over 16 runs for each set of parameters. For the tile-imprinting we changed randomly the wave vector directions within the tiles, for pixel-imprinting we change the pixel phases and for the spoon-stirring we slightly change the rotation orbit diameter.
@@ III.1 Comparison of imprinting techniques and spoon-stirring vortex injection
Fig. 2 shows the results of vortex detection and the application of the clustering algorithm for the cases of the Tile-8 phase-imprinting technique and for the spoon-stirring procedure at the beginning of the analysis phase (for conservative simulation). The more detailed simulation snapshot series and links to the videos are given in Appendix. For the spoon-stirring, the chosen simulation time was enough to generate significant amount of quantized vortices and their clusters. For the tile-imprinting, during the first several dozens of picoseconds, stabilization of vortices seeded by the initial phase singularities occurred. Further, the long-wavelength density waves extinction and quantum fluid homogenization took place. Visually, the resulting vortex distributions and their organization into clusters nearly coincide for Tile-8 and spoon-stirring. However, a detailed quantitative analysis of the excitation strategies is necessary from the point of view of cluster statistics and IKE spectra.
@@
To do so, we plot histograms to highlight vortex clustering for all excitation strategies employed. The left counterparts of the paired columns in panel a) in Fig. 3 show the number of single vortices, vortices in dipoles and vortices in clusters at the beginning of the analysis stage for all excitation techniques employed. One can see that Tile-8 and Tile-12 strategies resemble the spoon-stirring one much better than other strategies. Fractions of vortices in single state, dipoles and clusters are sufficiently similar. Due to this coincidence, Tile-8 with tile size of 32 m (see Fig. S3 for alternative grids) can be chosen as a reference phase-imprinting method for more detailed comparison with the spoon stirring approach. Setting decay parameter to zero during the preparatory part had rather quantitative than qualitative effect as it is shown in Appendix. For Tile-4 and Pixel strategies, one observes substantially lower fraction of vortices in clusters. The pixel schemes do not provide the global rotational motion and Tile-4 scheme injects too high kinetic energy with respect to potential one. Panel b) in Fig. 3 gives more details on clustering for spoon stirring and phase imprinting Tile-8 strategy. Tile-8 gives a slightly higher number of vortices in large clusters whereas the spoon-stirring strategy produces more single vortices, dipoles and small clusters. Typical intervortex distance can be estimated as , which gives m for Tile-8 and spoon-stirring strategies. The dominant fraction of clustered vortices is concentrated in large clusters.
@@
IKE spectra calculated via positions of all vortices and via positions of clustered vortices only (employing Eq. (2)) are depicted in panel a) of Fig. 4 for all excitation strategies. One can see very similar full IKE spectra regardless of the strategy employed; the discrepancies are within the values imposed by the difference in the total number of vortices, see Fig. 3. The full IKE spectra are composed of two parts different in characteristic power law. Exponent -3 corresponds to vortex cores at high wave vectors (greater than ) and it is presented in IKE spectra for all excitation methods and analysis frameworks. The exponent -1 dominates at lower wave vectors, being a fingerprint of an uncorrelated gas of vortices with dominating dipoles and single vortices. At the same time, the spectra for clustered vortices demonstrate a visible region of -5/3 power law, as was shown earlier for different excitation strategies [37]. This region spans from to , where the high-scale boundary is defined by a tile size as . The system has a certain memory effect of the tile size visible by the energy surplus stored at the wave vector , see Appendix, Fig. S4, S5. From to the slope -1 takes place. It should be noted that the Tile-8 phase-imprinting excitation strategy demonstrates full and clustered vortex-based IKE spectra very similar to those of the spoon-stirring approach. Among the phase-imprinting techniques, Tile-8 yields the highest clusterization degree comparable to the spoon stirring case. For pixel-type imprinting contribution from clustered vortices is lower, which agrees with the histograms. Panels b) and c) in Fig. 4 show full IKE spectra obtained semi-analytically via Eq. (2) (for full IKE spectra and IKE spectra of clustered vortices) and numerically via Eq. (6) for Tile-8 and spoon stirring cases, respectively. One sees that at low wave vectors (large size scales) full and clustered vortex IKE spectra coincide. This is due to the vanishing global rotational motion from the uncorrelated part of the “vortex gas”. For the wave vectors above , one sees that full IKE spectra have higher magnitude than the clustered vortex IKE spectra proportionally to the fraction of clustered vortices. The compressible energy contribution is lower than that of the incompressible, see Fig. S5.
@@ III.2 Effects of losses on sustainability of quantum turbulence
In present section, we investigate at which extent quantum fluid density depletion effects on IKE spectra and vortex statistics of quantum fluid with losses (radiative losses in case of polaritons). At this stage of the analysis, it has been clear that Tile-8 imprinting is a valid quantum fluid excitation method to study quantum turbulence, as shown by the analysis of the IKE spectra and vortex statistics. So, we focus on this excitation method. First, we track the time evolution of quantum fluid density and total number of vortices, see Fig. 5 panels a) and b), respectively. One sees that for high lifetime the decay of the total number of vortices is very close to the conservative case. For stronger decay, the decrease of the total number of vortices is stronger. When the density reaches approximately 20 (or 10% of density at the beginning of the analysis phase), an abrupt increase in the number of vortices in the automatized detection procedure becomes visible, which corresponds to turbulence vanishment associated with density depletion and appearance of the noise in the phase field, see Appendix Fig. S6. However, for lifetimes higher than 1 ps, turbulence exists for the whole observation period.
@@ III.3 Quantum turbulence in the regimes of fluctuating gain and loss
In the present section, we investigate the possibility of maintaining quantum turbulence in the regime with alternating phases of positive and negative gain. For polaritons, this regime approximates the system behavior under incoherent laser pump with time-dependent power operating above and below the threshold. We intend to characterize the effects of gain/loss fluctuations, possibly arising either from intentionally modulated lasers or from experimental limitations on the optical laser stability.